OpenStudy (anneauthor):

what is the arc length of the arc subtended in a circle with a radius of 6 and an angle of (7 * pi)/8

OpenStudy (khantahmina):

L=(7pi/8)(pi)(6)/(180) L=(157.5º)(pi)/30 L=(315º)(pi)/60 L=(21)(pi)/4 L=21pi/4

OpenStudy (khantahmina):

so i just found this here: what is the arc length of the arc subtended in a circle with radius 6 and ... https://www.wyzant.com › Resources › Answers